The Tyler family had heating bills of $105,$110,$90, and $67 in the first 4 months of 2003. the bills for the same months in 2004 wre $110,$95,$75, and $76. in which year was the mean monthly bill higher and by how much?
Answer by doukungfoo(195)

When asked to find the mean of a group of numbers we are being asked to find the average of that group. To find the average add them all together and divide by how many there are.
So for year 2003 we are given four heating bills. Lets add them together.
105 + 110 + 90 + 67 = 372
now divide 372 by 4
372/4 = 93
$93 is the mean for year 2003

Now lets find the mean for year 2004
again first add them all together
110 + 95 + 75 + 76 = 356
and then divide by 4
356/4 = 89
$89 is the mean for year 2004

In which year was the mean monthly bill higher and by how much?
year 2003 had the higher mean monthly bill by $4